# Non-perturbative determination of the sphaleron rate for first-order phase transitions

### v1

@@ Abstract
In many extensions of the Standard Model electroweak phase transitions at high temperatures can be described in a minimal dimensionally reduced effective theory with SU(2) gauge field and fundamental Higgs scalar. In this effective theory, all thermodynamic information is governed by two dimensionless ratios $x\equiv\lambda_{3}/g^{2}_{3}$ and $y\equiv m^{2}_{3}/g^{4}_{3}$, where $\lambda_{3}$, $m^{2}_{3}$ and $g_{3}$ are the effective thermal scalar self-interaction coupling, the thermal mass and the effective gauge-coupling, respectively. By using non-perturbative lattice simulations to determine the rate of sphaleron transitions in the entire $(x,y)$-plane corresponding to the Higgs phase, and by applying previous lattice results for the bubble nucleation, we find a condition $x(T_{c})\lesssim 0.025$ to guarantee preservation of the baryon asymmetry, which translates to $v/T_{c}\equiv\sqrt{2\langle\phi^{\dagger}\phi\rangle}/T_{c}\gtrsim 1.33$ for the (gauge-invariant) Higgs condensate. This indicates that viability of the electroweak baryogenesis requires the phase transition to be slightly stronger than previously anticipated. Finally, we present a general template for analysing such viability in a wide class of beyond the Standard Model theories, in which new fields are heavy enough to be integrated out at high temperature.
@@ I Introduction
Explanation to the observed matter/antimatter asymmetry of the universe still remains elusive, despite long-standing efforts Bodeker and Buchmuller (2021). Among the most studied candidates for a dynamical generation of the asymmetry is the electroweak baryogenesis Rubakov and Shaposhnikov (1996); Morrissey and Ramsey-Musolf (2012), where baryonic excess over their antimatter counterparts is generated during an electroweak phase transition. Such a phase transition is a violent process, where macroscopic bubbles of the low-temperature Higgs phase nucleate and eventually fill the young cosmos. Since the detection of gravitational waves from astrophysical sources a decade ago Abbott et al. (2016), attempts to understand such phase transitions have become fashionable, due to their potential to source a stochastic background of gravitational waves Caprini et al. (2016, 2020); Auclair et al. (2023). Detecting such a background is among the objectives of the future LISA mission Amaro-Seoane et al. (2017) and other next-generation space-borne gravitational wave experiments Kawamura et al. (2006); Harry et al. (2006); Ruan et al. (2020).
While the Standard Model of particle physics (SM) with the minimal scalar sector lacks a first-order electroweak phase transition Kajantie et al. (1996a); Aoki (1997); Gurtler et al. (1997); Csikor et al. (1999); Rummukainen et al. (1998), in many of its extensions a first-order transition can be triggered Cline and Kainulainen (1996); Laine (1996); Laine and Rummukainen (1998a, b, 2001); Laine et al. (2013); Kainulainen et al. (2019); Niemi et al. (2021a, 2024). This requires that new scalar fields, relevant at the electroweak scale, are light enough and are not too weakly interacting with the Higgs field, making such models attractive targets for future collider experiments Ramsey-Musolf (2020).
@@
While such a process has multiple stages, in this article our goal is to tackle only one of them: to accurately determine the sphaleron rate in scenarios that have a first-order electroweak phase transition. To achieve this, we resort to non-perturbative lattice simulations in a minimal dimensionally reduced effective theory for the electroweak phase transition. Such a minimal setup is provided by the theory with SU(2) gauge field and fundamental Higgs scalar, for which the ratio of thermal effective Higgs self-interaction coupling and the effective gauge-coupling $x\equiv\lambda_{3}/g^{2}_{3}$ at the critical temperature controls the character of phase transition, first-order transitions existing for $x<x_{*}=0.0983$ Kajantie et al. (1996b); Gould et al. (2022). This condition can be fulfilled in a class of theories where new scalars are sufficiently heavy at high temperatures (but not necessarily at zero temperature) that they can be integrated out from the thermal effective theory, while reducing the effective Higgs self-interaction coupling and thereby rendering the phase transition to be of first-order. Typically, this requires a large portal interaction with the Higgs field, enhancing detection prospects with future particle colliders.
@@
Furthermore, as a by-product, we also provide a comprehensive comparison to a recent fully perturbative approach tackling the same problem Li et al. (2025). Such comparison has also further-reaching motives: compared to perturbation theory, non-perturbative lattice studies are expensive, taking considerable computation time. At present, such simulations for the sphaleron rate are well established for the minimal effective theory we work with Moore (1999, 2000a); D’Onofrio et al. (2012, 2014a); Annala and Rummukainen (2023), as well as for the bubble nucleation rate Moore and Rummukainen (2001); Gould et al. (2022) (see also Moore et al. (2001); Gould et al. (2025); Hällfors and Rummukainen (2025)). Yet, for more complicated models, which include new fields within the effective theory, similar non-perturbative studies have not yet been performed, albeit non-perturbative simulations of equilibrium properties have appeared in Laine and Rummukainen (1998a, b, 2001); Laine et al. (2013); Kainulainen et al. (2019); Niemi et al. (2021a, 2024). However, establishing a reliable perturbative approach – i.e. concretely learning which perturbative effects have to be included and what are their limitations – makes it possible to tackle these more complicated models by purely perturbative approach, which would allow a scan through their often vast parameter spaces with relative ease.
@@ II Setup for the simulations
The sphaleron rate111 The term sphaleron rate is often used loosely to refer to the Chern-Simons number diffusion constant in both phases, while technically the term sphaleron coined in Klinkhamer and Manton (1984) refers to the classical saddle-point configuration, which only exists in the Higgs phase. Also, one often encounters alternative terminology for the two phases “symmetric/broken”, yet we avoid this as local gauge symmetry cannot be spontaneously broken Elitzur (1975). is the diffusion constant for the Chern-Simons number $N_{\rm CS}$, which is defined through the SU(2) field strength $F_{\alpha\beta}$ as Kuzmin et al. (1985); Rubakov and Shaposhnikov (1996); Moore and Tassler (2011)
@@
Akin to Farakos et al. (1995); D’Onofrio et al. (2012, 2014a); Annala and Rummukainen (2023) our starting point for the simulations is a dimensionally reduced effective field theory (EFT) in three spatial dimensions (3D) Ginsparg (1980); Appelquist and Pisarski (1981); Braaten and Nieto (1995); Kajantie et al. (1996c), defined through the following continuum action [^2] In this EFT, temporal gauge field components have been integrated out Kajantie et al. (1996c). For validity of this, see Appendix A.
@@
via thermal scalar self-interaction and thermal mass, respectively. We do not include U(1) gauge field (i.e. we set the Weinberg angle $\theta_{W}=0$) within the EFT, since it has a negligible effect on the phase transition Kajantie et al. (1996a, 1997); D’Onofrio and Rummukainen (2016); Laine and Meyer (2015) and the sphaleron rate333 We assume here that there are no strong external hypermagnetic fields. Strong hypermagnetic fields affect both the phase transition and the sphaleron rate, see e.g. refs. Annala and Rummukainen (2023); Kajantie et al. (1999) for non-perturbative studies. Klinkhamer and Laterveer (1992); Kleihaus et al. (1991); Kunz et al. (1992); Annala and Rummukainen (2023). Eq. (2) is the effective theory of the SM, but also of many of its extensions, provided that new BSM fields are heavy enough (akin to temporal gauge field components) at and below the critical temperature of the transition. We provide concrete examples of this in Sec. V.
Phase diagram of this EFT is dictated by the curve $y(T_{c})=y_{c}(x)$, which has been determined non-perturbatively in Kajantie et al. (1996b); Gould et al. (2022) (as well as perturbatively in Kajantie et al. (1996b); Ekstedt et al. (2022, 2024) to maximal perturbatively available order Linde (1980)). Below the critical $y_{c}(x)$ curve (after supercooling to temperatures below the critical temperature), one can associate a similar curve $y_{p}(x)$ for the condition that the system percolates, and the transition completes Gould et al. (2022). This curve can be determined from combining a condition from cosmological evolution with the bubble nucleation rate, which has been measured non-perturbatively for four different values of $x$ in Gould et al. (2022), as well as calculated perturbatively in Baacke and Heitmann (1999); Ekstedt (2022a, b); Gould et al. (2022).
Here, we find non-perturbatively an analogous curve $y_{f}(x)$ for a condition of the sphaleron freeze-out in the Higgs phase Burnier et al. (2006). Recently, such a curve was determined in a semi-classical approximation within the EFT in Li et al. (2025). As formulated in Li et al. (2025), if the sphaleron freeze-out occurs at temperatures higher than the percolation temperature, any baryon number washout is prevented inside the bubble, and hence any excess generated at the bubble wall would automatically be preserved inside the bubble, potentially leading to successful baryogenesis.444 We note, that this conclusion relies on an assumption that the system remains in thermal equilibrium, yet across the bubble wall the system departs from equilibrium and temperature varies across the bubble. For deflagrations, the bubble can heat the plasma in the Higgs phase inside the bubble back towards the critical temperature, c.f Enqvist et al. (1992). We have implicitly assumed that such out-of-equilibrium effects are negligible.
@@
Following previous works D’Onofrio et al. (2012, 2014a); Annala and Rummukainen (2023) we use the effective Langevin type description for the dynamical evolution. The dynamics of the soft modes that govern the evolution of the sphaleron configurations is fully overdamped and given by an effective Langevin type evolution to a leading logarithmic accuracy $1/\ln(1/g)$ Bodeker (1998); Arnold et al. (1999a); Bodeker (1999); Arnold et al. (1999b); Arnold and Yaffe (2000) in terms of SU(2) coupling $g$ of the parent theory.
In the Higgs phase the sphaleron rate can be computed from the lattice as Moore (1999); Moore and Rummukainen (2001)555 Technically, this is the Chern-Simons number diffusion rate, which is twice the sphaleron rate Burnier et al. (2006). However, we refer Eq. (8) as the sphaleron rate in order to stay consistent with terminology used in Annala and Rummukainen (2023).
@@
where $m_{D}^{2}$ is the Debye mass. In the SM $m_{D}^{2}=(11/6)g^{2}T^{2}$ leading to the numerical value $\sigma_{\text{el}}=0.91895T$. In principle, the BSM physics would effect $\sigma_{\textrm{el}}$ through modifying the matching relation for the Debye mass, as well as hard thermal loops used to derive Eq. (II), c.f. Moore (2000b, a). Such effects can be expected to have minor importance, and for simplicity we do not consider them further.
@@ IV Cosmology: The Sphaleron Freeze-out
Figure 4: The sphaleron rate in the $(x,y)$ plane. The contour and black dotted constant rate lines are obtained from linearly interpolating the data obtained from simulations with $ag_{3}^{2}=0.25$, $V=32^{3}a^{3}$. The orange almost vertical line corresponds to the SM mapping for temperatures $130-167~{}$GeV. The gray dots with error bars are the sphaleron freeze-out $y$-values for fixed $x$, i.e. $y_{f}(x)$. The black solid line is a fit to small $x\lesssim 0.03$ values of $y_{f}$. The red and blue solid lines are the non-perturbative results for $y_{c}(x)$ and $y_{p}(x)$ respectively Gould et al. (2022). The star denotes the first-order phase transition end point $x_{*}=0.0983$ Gould et al. (2022), i.e. for larger values of $x$ the transition turns into a smooth crossover. On the left panel the dot-dashed box denotes the area that is zoomed in on the right panel which shows the relevant range for strong first-order phase transitions.
@@
where $N_{\rho}\approx\frac{13}{4}\times 3$, with the Hubble parameter $H(T)=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{T^{2}}{m_{\text{pl}}}$, with $g_{*}$ relativistic degrees of freedom and the Planck mass $m_{\text{pl}}=1.22\times 10^{19}$ GeV. From Eq. (11) we can solve $T_{f}$ for any dimensional reduction mapping exactly, but in analogy to definition of $y_{p}$ curve, we can get an approximate general result by noting that
@@
is approximately a constant for the relevant temperature range. For example, for $T=100$ to 250250250~{}250GeV right-hand side of (12) varies only from $-38.8$ to $-37.9$. Furthermore, we assume that $g^{2}_{3}$ and $\sigma_{\textrm{el}}$ are not significantly affected by the BSM physics. This allows us to invert this condition, to find a curve $y_{f}(x)$.
For a fixed value of $x$ we are not able to efficiently simulate $y$ values that are very deep in the Higgs phase far from the critical $y_{c}$ value, and thus for larger $x$ values ($x\gtrsim 0.1$) we cannot reach the point of the freeze-out condition. However, the rate in the Higgs phase is well approximated by a pure exponential, and thus we can perform a linear fit to $\ln\tilde{\Gamma}(y)$ to extrapolate the rate deeper to the Higgs phase to find the $y_{f}$ freeze-out value (similarly what was done in the SM case D’Onofrio et al. (2014b); Annala and Rummukainen (2023)), see e.g. Fig. 3. Regardless, for the interesting region of first-order transitions with small $x\lesssim 0.1$ we do not need to extrapolate to deeper in the Higgs phase.
@@
Figure 5: Zoomed in plot of Fig. 4 where the sphaleron freeze-out $y_{f}$ and the nucleation temperature $y_{p}$ curves intersect at $0.0251^{+0.0012}_{-0.0002}$. The error bounds for the blue $y_{p}$ curve are from varying the cosmology condition ±$\pm 5$. The black solid line is a linear fit to the obtained $y_{f}(x)$ values for $x<0.03$ and the gray error bound is a combination of the fit error and the estimated continuum limit error. The linear fit for small $x$ is $y_{f}=(-5.65\pm 0.14)x+(0.164\pm 0.004)$.
@@
prevent the sphaleron washout inside the bubbles of the Higgs phase, and hence $x_{c}\lesssim\bar{x}$ guarantees baryon number preservation after strong first-order phase transition. We point out, that this condition is more strict compared to conditions $x_{c}\lesssim 0.04$ reported in the seminal work Kajantie et al. (1996b) as well as $x_{c}\lesssim 0.037$ obtained in Moore (1999). This indicates, that for viable baryogenesis the electroweak phase transition has to be somewhat stronger than previously anticipated.

### v2

@@ Abstract
In many extensions of the Standard Model electroweak phase transitions at high temperatures can be described in a minimal dimensionally reduced effective theory with SU(2) gauge field and fundamental Higgs scalar. In this effective theory, all thermodynamic information is governed by two dimensionless ratios $x\equiv\lambda_{3}/g^{2}_{3}$ and $y\equiv m^{2}_{3}/g^{4}_{3}$, where $\lambda_{3}$, $m^{2}_{3}$ and $g_{3}$ are the effective thermal scalar self-interaction coupling, the thermal mass and the effective gauge-coupling, respectively. By using non-perturbative lattice simulations to determine the rate of sphaleron transitions in the entire $(x,y)$-plane corresponding to the Higgs phase, and by applying previous lattice results for the bubble nucleation, we find a condition $x(T_{c})\lesssim 0.025$ to guarantee preservation of the baryon asymmetry, which translates to $v/T_{c}\equiv\sqrt{2\Delta\langle\phi^{\dagger}\phi\rangle}/T_{c}\gtrsim 1.33$ for the (gauge-invariant) discontinuity in Higgs condensate. This indicates that viability of the electroweak baryogenesis requires the phase transition to be slightly stronger than previously anticipated. Finally, we present a general template for analysing such viability in a wide class of beyond the Standard Model theories, in which new fields are heavy enough to be integrated out at high temperature.
@@ I Introduction
Explanation to the observed matter/antimatter asymmetry of the universe still remains elusive, despite long-standing efforts Bodeker and Buchmuller (2021). Among the most studied candidates for a dynamical generation of the asymmetry is the electroweak baryogenesis Rubakov and Shaposhnikov (1996); Morrissey and Ramsey-Musolf (2012), where baryonic excess over their antimatter counterparts is generated during an electroweak phase transition. Such a phase transition is a violent process, where macroscopic bubbles of the low-temperature Higgs phase nucleate and eventually fill the young cosmos. Since the detection of gravitational waves from astrophysical sources a decade ago Abbott and others (2016), attempts to understand such phase transitions have become fashionable, due to their potential to source a stochastic background of gravitational waves Caprini and others (2016, 2020); Auclair and others (2023). Detecting such a background is among the objectives of the future LISA mission Amaro-Seoane and others (2017) and other next-generation space-borne gravitational wave experiments Kawamura and others (2006); Harry et al. (2006); Ruan et al. (2020).
While the Standard Model of particle physics (SM) with the minimal scalar sector lacks a first-order electroweak phase transition Kajantie et al. (1996b); Aoki (1997); Gurtler et al. (1997); Csikor et al. (1999); Rummukainen et al. (1998), in many of its extensions a first-order transition can be triggered Cline and Kainulainen (1996); Laine (1996); Laine and Rummukainen (1998b, a, 2001); Laine et al. (2013); Kainulainen et al. (2019); Niemi et al. (2021a, 2024). This requires that new scalar fields, relevant at the electroweak scale, are light enough and are not too weakly interacting with the Higgs field, making such models attractive targets for future collider experiments Ramsey-Musolf (2020).
@@
While such a process has multiple stages, in this article our goal is to tackle only one of them: to accurately determine the sphaleron rate in scenarios that have a first-order electroweak phase transition. To achieve this, we resort to non-perturbative lattice simulations in a minimal dimensionally reduced effective theory for the electroweak phase transition. Such a minimal setup is provided by the theory with SU(2) gauge field and fundamental Higgs scalar, for which the ratio of thermal effective Higgs self-interaction coupling and the effective gauge-coupling $x\equiv\lambda_{3}/g^{2}_{3}$ at the critical temperature controls the character of phase transition, first-order transitions existing for $x<x_{*}=0.0983$ Kajantie et al. (1996c); Gould et al. (2022). This condition can be fulfilled in a class of theories where new scalars are sufficiently heavy at high temperatures (but not necessarily at zero temperature) that they can be integrated out from the thermal effective theory, while reducing the effective Higgs self-interaction coupling and thereby rendering the phase transition to be of first-order. Typically, this requires a large portal interaction with the Higgs field, enhancing detection prospects with future particle colliders.
@@
Furthermore, as a by-product, we also provide a comprehensive comparison to a recent fully perturbative approach tackling the same problem Li et al. (2025). Such comparison has also further-reaching motives: compared to perturbation theory, non-perturbative lattice studies are expensive, taking considerable computation time. At present, such simulations for the sphaleron rate are well established for the minimal effective theory we work with Moore (1999, 2000a); D’Onofrio et al. (2012, 2014b); Annala and Rummukainen (2023), as well as for the bubble nucleation rate Moore and Rummukainen (2001); Gould et al. (2022) (see also Moore et al. (2001); Gould et al. (2025); Hällfors and Rummukainen (2025)). Yet, for more complicated models, which include new fields within the effective theory, similar non-perturbative studies have not yet been performed, albeit non-perturbative simulations of equilibrium properties have appeared in Laine and Rummukainen (1998b, a, 2001); Laine et al. (2013); Kainulainen et al. (2019); Niemi et al. (2021a, 2024). However, establishing a reliable perturbative approach – i.e. concretely learning which perturbative effects have to be included and what are their limitations – makes it possible to tackle these more complicated models by purely perturbative approach, which would allow a scan through their often vast parameter spaces with relative ease.
@@ II Setup for the simulations
The sphaleron rate [^1] The term sphaleron rate is often used loosely to refer to the Chern-Simons number diffusion constant in both phases, while technically the term sphaleron coined in Klinkhamer and Manton (1984) refers to the classical saddle-point configuration, which only exists in the Higgs phase. Also, one often encounters alternative terminology for the two phases “symmetric/broken”, yet we avoid this as local gauge symmetry cannot be spontaneously broken Elitzur (1975). is the diffusion constant for the Chern-Simons number $N_{\rm CS}$, which is defined through the SU(2) field strength $F_{\alpha\beta}$ as Kuzmin et al. (1985); Rubakov and Shaposhnikov (1996); Moore and Tassler (2011)
@@
Akin to Farakos et al. (1995); D’Onofrio et al. (2012, 2014b); Annala and Rummukainen (2023) our starting point for the simulations is a dimensionally reduced effective field theory (EFT) in three spatial dimensions (3D) Ginsparg (1980); Appelquist and Pisarski (1981); Braaten and Nieto (1995); Kajantie et al. (1996a), defined through the following continuum action [^2] In this EFT, temporal gauge field components have been integrated out Kajantie et al. (1996a). For validity of this, see Appendix A.
@@
via thermal scalar self-interaction and thermal mass, respectively. We do not include U(1) gauge field (i.e. we set the Weinberg angle $\theta_{W}=0$) within the EFT, since it has a negligible effect on the phase transition Kajantie et al. (1996b, 1997); D’Onofrio and Rummukainen (2016); Laine and Meyer (2015) and the sphaleron rate [^3] We assume here that there are no strong external hypermagnetic fields. Strong hypermagnetic fields affect both the phase transition and the sphaleron rate, see e.g. refs. Annala and Rummukainen (2023); Kajantie et al. (1999) for non-perturbative studies. Klinkhamer and Laterveer (1992); Kleihaus et al. (1991); Kunz et al. (1992); Annala and Rummukainen (2023). Eq. (2) is the effective theory of the SM, but also of many of its extensions, provided that new BSM fields are heavy enough (akin to temporal gauge field components) at and below the critical temperature of the transition. We provide concrete examples of this in Sec. V.
Phase diagram of this EFT is dictated by the curve $y(T_{c})=y_{c}(x)$, which has been determined non-perturbatively in Kajantie et al. (1996c); Gould et al. (2022) (as well as perturbatively in Kajantie et al. (1996c); Ekstedt et al. (2022, 2024) to maximal perturbatively available order Linde (1980)). Below the critical $y_{c}(x)$ curve (after supercooling to temperatures below the critical temperature), one can associate a similar curve $y_{p}(x)$ for the condition that the system percolates, and the transition completes Gould et al. (2022). This curve can be determined from combining a condition from cosmological evolution with the bubble nucleation rate, which has been measured non-perturbatively for four different values of $x$ in Gould et al. (2022), as well as calculated perturbatively in Baacke and Heitmann (1999); Ekstedt (2022b, a); Gould et al. (2022).
Here, we find non-perturbatively an analogous curve $y_{f}(x)$ for a condition of the sphaleron freeze-out in the Higgs phase Burnier et al. (2006). Recently, such a curve was determined in a semi-classical approximation within the EFT in Li et al. (2025). As formulated in Li et al. (2025), if the sphaleron freeze-out occurs at temperatures higher than the percolation temperature, any baryon number washout is prevented inside the bubble, and hence any excess generated at the bubble wall would automatically be preserved inside the bubble, potentially leading to successful baryogenesis. [^4] We note, that this conclusion relies on an assumption that the system remains in thermal equilibrium, yet across the bubble wall the system departs from equilibrium and temperature varies across the bubble. For deflagrations, the bubble can heat the plasma in the Higgs phase inside the bubble back towards the critical temperature, c.f Enqvist et al. (1992). We have implicitly assumed that such out-of-equilibrium effects are negligible.
@@
Following previous works D’Onofrio et al. (2012, 2014b); Annala and Rummukainen (2023) we use the effective Langevin type description for the dynamical evolution. The dynamics of the soft modes that govern the evolution of the sphaleron configurations is fully overdamped and given by an effective Langevin type evolution to a leading logarithmic accuracy $1/\ln(1/g)$ Bodeker (1998); Arnold et al. (1999a); Bodeker (1999); Arnold et al. (1999b); Arnold and Yaffe (2000) in terms of SU(2) coupling $g$ of the parent theory.
In the Higgs phase the sphaleron rate can be computed from the lattice as Moore (1999); Moore and Rummukainen (2001) [^5] Technically, this is the Chern-Simons number diffusion rate, which is twice the sphaleron rate Burnier et al. (2006). However, we refer Eq. (8) as the sphaleron rate in order to stay consistent with terminology used in Annala and Rummukainen (2023).
@@
where $m_{D}^{2}$ is the Debye mass. In the SM $m_{D}^{2}=(11/6)g^{2}T^{2}$ leading to the numerical value $\sigma_{\text{el}}=0.91895T$ (see Bödeker and Klose (2025) for a recently computed QCD correction). In principle, the BSM physics would effect $\sigma_{\textrm{el}}$ through modifying the matching relation for the Debye mass, as well as hard thermal loops used to derive Eq. (II), c.f. Moore (2000b, a). Such effects can be expected to have minor importance, and for simplicity we do not consider them further.
@@ IV Cosmology: The Sphaleron Freeze-out
Figure 4: The sphaleron rate in the $(x,y)$ plane. The contour and black dotted constant rate lines are obtained from linearly interpolating the data obtained from simulations with $ag_{3}^{2}=0.25$, $V=32^{3}a^{3}$. The orange almost vertical line corresponds to the SM mapping for temperatures 130167130-167~GeV. The gray dots with error bars are the sphaleron freeze-out $y$-values for fixed $x$, i.e. $y_{f}(x)$. The black solid line is a fit to small $x\lesssim 0.03$ values of $y_{f}$. The red and blue solid lines are the non-perturbative results for $y_{c}(x)$ and $y_{p}(x)$ respectively Gould et al. (2022). The star denotes the first-order phase transition end point $x_{*}=0.0983$ Gould et al. (2022), i.e. for larger values of $x$ the transition turns into a smooth crossover. On the left panel the dot-dashed box denotes the area that is zoomed in on the right panel which shows the relevant range for strong first-order phase transitions.
@@
[…]
@@
is approximately a constant for the relevant temperature range. For example, for $T=100$ to 250250~GeV right-hand side of (12) varies only from 38.8-38.8 to 37.9-37.9. Furthermore, we assume that $g^{2}_{3}$ and $\sigma_{\textrm{el}}$ are not significantly affected by the BSM physics. This allows us to invert this condition, to find a curve $y_{f}(x)$.
For a fixed value of $x$ we are not able to efficiently simulate $y$ values that are very deep in the Higgs phase far from the critical $y_{c}$ value, and thus for larger $x$ values ($x\gtrsim 0.1$) we cannot reach the point of the freeze-out condition. However, the rate in the Higgs phase is well approximated by a pure exponential, and thus we can perform a linear fit to $\ln\tilde{\Gamma}(y)$ to extrapolate the rate deeper to the Higgs phase to find the $y_{f}$ freeze-out value (similarly what was done in the SM case D’Onofrio et al. (2014a); Annala and Rummukainen (2023)), see e.g. Fig. 3. Regardless, for the interesting region of first-order transitions with small $x\lesssim 0.1$ we do not need to extrapolate to deeper in the Higgs phase.
@@
Figure 5: Zoomed in plot of Fig. 4 where the sphaleron freeze-out $y_{f}$ and the nucleation temperature $y_{p}$ curves intersect at $0.0251^{+0.0012}_{-0.0002}$. The error bounds for the blue $y_{p}$ curve are from varying the cosmology condition $\pm 5$. The black solid line is a linear fit to the obtained $y_{f}(x)$ values for $x<0.03$ and the gray error bound is a combination of the fit error and the estimated continuum limit error. The linear fit for small $x$ is $y_{f}=(-5.65\pm 0.14)x+(0.164\pm 0.004)$.
@@
prevent the sphaleron washout inside the bubbles of the Higgs phase, and hence $x_{c}\lesssim\bar{x}$ guarantees baryon number preservation after strong first-order phase transition. We point out, that this condition is more strict compared to conditions $x_{c}\lesssim 0.04$ reported in the seminal work Kajantie et al. (1996c) as well as $x_{c}\lesssim 0.037$ obtained in Moore (1999). This indicates, that for viable baryogenesis the electroweak phase transition has to be somewhat stronger than previously anticipated.